OpenStudy (anonymous):

The parking lot behind Kai's Kafe has a perimeter of 210 m. The width is three-fourths of the length. What are the dimensions of the parking lot?

OpenStudy (anonymous):

P(L)= (7/2)L

OpenStudy (anonymous):

Generally, P=2L+2w for rectangles with length L and width w. w=3/4*L, so we substitute it in for w. 210=2L+2(3/4*L) Now, we solve. 210=2L+6/4L 210=4/2L+3/2L 210=7/2L ***60=L*** w=3/4L w=3/4*60 ***w=45*** The parking lot is 60m by 45m.